# TemPrompt: Multi-Task Prompt Learning for Temporal Relation Extraction in RAG-based Crowdsourcing Systems

### v2

@@ Abstract
Temporal relation extraction (TRE) aims to grasp the evolution of events or actions, and thus shape the workflow of associated tasks, so it holds promise in helping understand task requests initiated by requesters in crowdsourcing systems. However, existing methods still struggle with limited and unevenly distributed annotated data. Therefore, inspired by the abundant global knowledge stored within pre-trained language models (PLMs), we propose a multi-task prompt learning framework for TRE (TemPrompt), incorporating prompt tuning and contrastive learning to tackle these issues. To elicit more effective prompts for PLMs, we introduce a task-oriented prompt construction approach that thoroughly takes the myriad factors of TRE into consideration for automatic prompt generation. In addition, we present temporal event reasoning as a supplement to bolster the model’s focus on events and temporal cues. The experimental results demonstrate that TemPrompt outperforms all compared baselines across the majority of metrics under both standard and few-shot settings. A case study is provided to validate its effectiveness in crowdsourcing scenarios.
Index Terms:
@@ I Introduction
  • We present an automatic generation approach for discrete prompts of TRE, which takes triggers, labels, and event mentions into consideration and requires no significant supervision costs.
@@ II Related Work
II-A Temporal Relation Extraction
@@ IV Model Architecture
IV-A Prompt-based Classifier
@@
To achieve more accurate predictions, the supervised samples $\{(S_{i},r_{i})\}$ are constructed and utilized to fine-tune the PLM $\mathcal{M}$ by minimizing the cross-entropy loss $\mathcal{L}_{ce}$:
@@
IV-B Task-oriented Prompt Construction
@@
IV-C Temporal Event Reasoning
@@ IV-E The Total Loss
V Experiments
V-A Datasets
We conduct a series of experiments on two public benchmark datasets for TRE: TB-Dense [53] and MATRES[54]. TB-Dense is based on TimeBank Corpus [55] and annotated densely with six types of relations: BEFORE, AFTER, SIMULTANEOUS, INCLUDES, INCLUDED IN and VAGUE. It requires annotating all event/time pairs within a specified window and introduces a VAGUE label to capture difficult-to-distinguish relations even by humans. MATRES further simplifies temporal relations into four distinct categories: BEFORE, AFTER, EQUAL and VAGUE. It offers refined TRE annotations on TimeBank [55] and TempEval [56] (including AQUAINT and Platinum subsets) documents. Our experimental setup for both benchmark datasets involves utilizing the same training/validation/testing splits as those used in previous studies [5, 40, 15]. Their detailed statistics and data split can be found in Table I.
TABLE I: Statistics of TB-Dense and MATRES datasets.
@@ V-B Baselines
V-C Training Settings and Evaluation Metrics
@@
TABLE II: Results of various TRE methods on MATRES.
@@
TABLE III: Results of various TRE methods on TB-Dense.
@@
V-D Results and analysis
The results of our models and various baselines on the MATRES and TB-Dense datasets are shown in Table II and III, with the best results highlighted in bold and the second-best results underlined. Except for the R score on MATRES, our approach outperforms all baseline models on the two datasets. Specifically, our TemPrompt (TRE) model achieves improvements of 0.3% and 3% in F1 scores on MATRES and TB-Dense, respectively, compared to state-of-the-art models. This demonstrates the effectiveness and superiority of our method in extracting temporal relations.
@@ V-D Results and analysis
TABLE IV: Ablation study on MATRES and TB-Dense.
@@
V-E Ablation Study
To evaluate the impact of each component on the overall performance of the TemPrompt model, we conduct ablation experiments separately on two datasets. We remove temporal event reasoning and contrastive learning individually, denoted as -wo TRE and -wo CL, for testing purposes. The results of ablation studies are shown in Table IV. After removing both, we observe a decline in metrics, highlighting the significant positive influence of both on the performance of TRE. However, the extent of performance decline varies across different datasets. The removal of temporal event reasoning leads to a decrease of 0.91% in F1 score on MATRES and a decrease of 0.7% on TB-Dense, while the absence of contrastive learning resulted in a 1.31 % decrease in F1 score on MATRES and a 0.31% decrease on TB-Dense. This might be attributed to the fact that in TB-Dense, two events unfold within continuous lengthy texts, yet we only merge sentences containing these two events. This disruption of contextual semantic continuity may impede overall contextual comprehension and consequently weaken the promotional effect of temporal event reasoning.
TABLE V: Results of different templates on MATRES.
@@ V-E Ablation Study
V-F Few-shot Event Study
@@
V-G Analysis of Different Templates
To assess the interpretability and effectiveness of various prompts, we examine the top five generated templates (i.e., Prompt1, Prompt2, Prompt3, Prompt4 and Prompt5) that demonstrate the best performance by TemPrompt on the small-scale dataset, as well as manually designed templates (i.e., Prompt0). Based on these templates, we evaluate and compare their results on the overall test set of MATRES, as shown in Table V. It is observed that the automatically generated templates exhibit grammatical accuracy, and coherent phrasing, and are easily interpretable. Additionally, they achieve higher F1 scores than manual templates. This can be attributed to the tendency for manually created templates to be suboptimal when attempting to elicit the correct response from the language model in tasks involving multiple categories. The model’s performance is sensitive to template selection, as illustrated by the 3.31% higher F1 score achieved by Prompt5 compared to Prompt1, despite both being automated templates. Therefore, our few-shot evaluation strategy is important and indispensable. This further exemplifies the superiority of our task-oriented prompt generation. Despite their apparent interpretability to humans, manually crafted templates often fail to align with the language model’s interpretation of inputs, thereby compromising model performance.
V-H Visualization of Predicted Representations
@@
V-I Label-wise Analysis
For a detailed analysis of the improvements made by our model in addressing data issues, we compare its performance with PSL on individual labels, as shown in Table VI. For labels with comparatively more samples, TemPrompt achieves improvements of 1.2%, 11.7% and 5.1% in F1 scores for the VAGUE, BEFORE and AFTER labels, respectively. In terms of labels with relatively fewer samples, despite a slight decrease for the INCLUDED IN label, TemPrompt demonstrates an 8.3% increase in F1 score compared to PSL for the INCLUDES label. Especially, concerning the SIMULTANEOUS label, referencing recent works [41, 34, 43], TemPrompt achieves a remarkable advancement that raises the F1 scores from 0 to 34.1%. Obviously, on the whole, our approach can effectively addresses the deficiencies and unevenness within human annotated data. However, it falls short of resolving all few-shot label challenges, highlighting the importance of further research.
TABLE VI: F1 scores for each individual label on TB-Dense.
@@ V-J Hyperparameter Analysis
V-K Case Study
To explore the application of our approach in RAG-based crowdsourcing, we input sentences about PCB design and manufacturing, which is the outcome retrieved by the retriever in response to the queries posted by the requester, into the fine-tuned TemPrompt on TB-Dense. Note that the events mentioned in the sentence have been detected in advance. Table VII presents the examples of TRE in the printed circuit board (PCB) manufacturing domain. Despite the lack of PCB-related data in the dataset, the model demonstrates its ability to effectively extract salient relations, i.e., BEFORE and AFTER. Nevertheless, the temporal relation between layout and creation is classified as BEFORE, but conversely, it is erroneously categorized as SIMULTANEOUS. This indicates one of the limitations of our model, prompting us to explore the integration of antisymmetric constraints in future research. For complex relations with subtle features, such as SIMULTANEOUS, VAGUE, INCLUDES and INCLUDED IN, we believe learning domain-specific knowledge is essential. For example, experts consider selection and layout to be in a continuous cycle of joint optimization, hence their temporal relation is annotated as SIMULTANEOUS. Additionally, they view procurement as an integral part of the manufacturing process, so the relation between the former and the latter is INCLUDED IN and conversely INCLUDES. Obviously, these mistakes stem from a lack of expertise in the specific field. In our future work, we plan to collect domain-specific knowledge or data so that the retriever can access it and improve our model.
TABLE VII: Examples of TRE in specialized domains.
@@ V-K Case Study
VI Conclusion and Future Work

### v4

@@ Abstract
Temporal relation extraction (TRE) aims to grasp the evolution of events or actions, and thus shape the workflow of associated tasks, so it holds promise in helping understand task requests initiated by requesters in crowdsourcing systems. However, existing methods still struggle with limited and unevenly distributed annotated data. Therefore, inspired by the abundant global knowledge stored within pre-trained language models (PLMs), we propose a multi-task prompt learning framework for TRE (TemPrompt), incorporating prompt tuning and contrastive learning to tackle these issues. To elicit more effective prompts for PLMs, we introduce a task-oriented prompt construction approach that thoroughly takes the myriad factors of TRE into consideration for automatic prompt generation. In addition, we design temporal event reasoning in the form of masked language modeling as auxiliary tasks to bolster the model’s focus on events and temporal cues. The experimental results demonstrate that TemPrompt outperforms all compared baselines across the majority of metrics under both standard and few-shot settings. A case study on designing and manufacturing printed circuit boards is provided to validate its effectiveness in crowdsourcing scenarios.
keywords:
@@ 1 Introduction
  • 2.
    We present a task-oriented automatic generation approach for discrete prompts of TRE, which takes triggers, labels, and event mentions into consideration and requires no significant supervision costs.
  • 3.
@@ 2 Related Work
2.1 Temporal Relation Extraction
@@ 4 Model Architecture
4.1 Prompt-based Classifier
@@
To achieve more accurate predictions, the supervised samples $\{(S_{i},r_{i})\}$ are constructed and utilized to fine-tune the PLM $\mathcal{M}$ by minimizing the cross-entropy loss $\mathcal{L}_{tre}$:
@@
4.2 Task-oriented Prompt Construction
@@
4.3 Temporal Event Reasoning
@@ 4.5 The Total Loss
5 Experiments
5.1 Datasets
We conduct a series of experiments on two public benchmark datasets for TRE: TB-Dense [53] and MATRES[54]. TB-Dense is based on TimeBank Corpus [55] and annotated densely with six types of relations: BEFORE, AFTER, SIMULTANEOUS, INCLUDES, INCLUDED IN and VAGUE. It requires annotating all event/time pairs within a specified window and introduces a VAGUE label to capture difficult-to-distinguish relations even by humans. MATRES further simplifies temporal relations into four distinct categories: BEFORE, AFTER, EQUAL and VAGUE. It offers refined TRE annotations on TimeBank [55] and TempEval [56] (including AQUAINT and Platinum subsets) documents. Our experimental setup for both benchmark datasets involves utilizing the same training/validation/testing splits as those used in previous studies [5, 40, 15]. Their detailed statistics and data split can be found in Table 1.
Table 1: Statistics of TB-Dense and MATRES datasets.
@@ 5.2 Baselines
5.3 Training Settings and Evaluation Metrics
@@
Table 2: Results of various TRE methods on MATRES.
@@
Table 3: Results of various TRE methods on TB-Dense.
@@
5.4 Results and analysis
The results of our models and various baselines on the MATRES and TB-Dense datasets are shown in Table 2 and 3, with the best results highlighted in bold and the second-best results underlined. Except for the R score on MATRES, our approach outperforms all baseline models on the two datasets. Specifically, our TemPrompt (TRE) model achieves improvements of 0.3% and 3% in F1 scores on MATRES and TB-Dense, respectively, compared to state-of-the-art models. This demonstrates the effectiveness and superiority of our method in extracting temporal relations.
@@ 5.4 Results and analysis
Table 4: Ablation study on MATRES and TB-Dense.
@@
5.5 Ablation Study
To evaluate the impact of each component on the overall performance of the TemPrompt model, we conduct ablation experiments separately on two datasets. We remove temporal event reasoning and contrastive learning individually, denoted as -wo TER and -wo CL, for testing purposes. The results of ablation studies are shown in Table 4. After removing both, we observe a decline in metrics, highlighting the significant positive influence of both on the performance of TRE. However, the extent of performance decline varies across different datasets. The removal of temporal event reasoning leads to a decrease of 0.91% in F1 score on MATRES and a decrease of 0.7% on TB-Dense, while the absence of contrastive learning resulted in a 1.31 % decrease in F1 score on MATRES and a 0.31% decrease on TB-Dense. This might be attributed to the fact that in TB-Dense, two events unfold within continuous lengthy texts, yet we only merge sentences containing these two events. This disruption of contextual semantic continuity may impede overall contextual comprehension and consequently weaken the promotional effect of temporal event reasoning.
Table 5: Results of different templates on MATRES.
@@ 5.5 Ablation Study
5.6 Few-shot Event Study
@@
5.7 Analysis of Different Templates
To assess the interpretability and effectiveness of various prompts, we examine the top five generated templates (i.e., Prompt1, Prompt2, Prompt3, Prompt4 and Prompt5) that demonstrate the best performance by TemPrompt on the small-scale dataset, as well as manually designed templates (i.e., Prompt0). Based on these templates, we evaluate and compare their results on the overall test set of MATRES, as shown in Table 5. It is observed that the automatically generated templates exhibit grammatical accuracy, and coherent phrasing, and are easily interpretable. Additionally, they achieve higher F1 scores than manual templates. This can be attributed to the tendency for manually created templates to be suboptimal when attempting to elicit the correct response from the language model in tasks involving multiple categories. The model’s performance is sensitive to template selection, as illustrated by the 3.31% higher F1 score achieved by Prompt5 compared to Prompt1, despite both being automated templates. Therefore, our few-shot evaluation strategy is important and indispensable. This further exemplifies the superiority of our task-oriented prompt generation. Despite their apparent interpretability to humans, manually crafted templates often fail to align with the language model’s interpretation of inputs, thereby compromising model performance.
5.8 Visualization of Predicted Representations
@@
5.9 Label-wise Analysis
For a detailed analysis of the improvements made by our model in addressing data issues, we compare its performance with PSL on individual labels, as shown in Table 6. For labels with comparatively more samples, TemPrompt achieves improvements of 1.2%, 11.7% and 5.1% in F1 scores for the VAGUE, BEFORE and AFTER labels, respectively. In terms of labels with relatively fewer samples, despite a slight decrease for the INCLUDED IN label, TemPrompt demonstrates an 8.3% increase in F1 score compared to PSL for the INCLUDES label. Especially, concerning the SIMULTANEOUS label, referencing recent works [41, 34, 43], TemPrompt achieves a remarkable advancement that raises the F1 scores from 0 to 34.1%. Obviously, on the whole, our approach can effectively addresses the deficiencies and unevenness within human annotated data. However, it falls short of resolving all few-shot label challenges, highlighting the importance of further research.
Table 6: F1 scores for each individual label on TB-Dense.
@@ 5.10 Hyperparameter Analysis
5.11 Case Study
To explore the application of our approach in RAG-based crowdsourcing, we input sentences about printed circuit board (PCB) design and manufacturing, which is the outcome retrieved by the retriever in response to the queries posted by the requester, into the fine-tuned TemPrompt on TB-Dense. Note that the events mentioned in the sentence have been detected in advance. Table 7 presents the examples of TRE in the PCB manufacturing domain. Despite the lack of PCB-related data in the dataset, the model demonstrates its ability to effectively extract salient relations, i.e., BEFORE and AFTER. Nevertheless, the temporal relation between layout and creation is classified as BEFORE, but conversely, it is erroneously categorized as SIMULTANEOUS. This indicates one of the limitations of our model, prompting us to explore the integration of antisymmetric constraints in future research. For complex relations with subtle features, such as SIMULTANEOUS, VAGUE, INCLUDES and INCLUDED IN, we believe learning domain-specific knowledge is essential. For example, experts consider selection and layout to be in a continuous cycle of joint optimization, hence their temporal relation is annotated as SIMULTANEOUS. Additionally, they view procurement as an integral part of the manufacturing process, so the relation between the former and the latter is INCLUDED IN and conversely INCLUDES. Obviously, these mistakes stem from a lack of expertise in the specific field. In our future work, we plan to collect domain-specific knowledge or data so that the retriever can access it and improve our model.
Table 7: Examples of TRE in specialized domains.
@@ 5.11 Case Study
6 Conclusion and Future Work
@@
CRediT authorship contribution statement
Jing Yang: Conceptualization, Methodology, Writing - Original Draft, Writing - Review & Editing. Yu Zhao: Methodology, Investigation, Validation. Linyao Yang: Writing - Review & Editing. Xiao Wang: Writing - Review & Editing, Funding acquisition. Long Chen: Writing - Review & Editing. Fei-Yue Wang: Writing - Review & Editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Acknowledgment
This work was supported in part by the Zhejiang Lab Open Research Project under Grant K2022KG0AB02, and in part by the University Scientific Research Program of Anhui Province under Grant 2023AH020005.